An aqueous solution of Borax (Na2B4O7) is sometimes used to fireproof wood. 


What is the molarity of 45.0 L of solution that contains 6.78 kg of Na2B4O7?


1
Expert's answer
2022-03-30T03:17:05-0400

CM = n / V

n = m / M

M (Na2B4O7) = 201.2 g/mol

n = 6780 / 201.2 = 33.7 mol

CM (Na2B4O7) = 33.7 / 45 = 0.75 M